# Robust Semiparametric Graphical Models with Skew-Elliptical Distributions

### v2

@@ 6.1 Numerical Simulations
Table 2: Quantitative comparisons of the 5 correlation matrix estimators using different distributions to generate data, and compute the $\Omega$ matrix. The graphs are estimated using the glasso algorithm. The values are in percentages.
SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) PEARSON
$n$ $p$ AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR
N DATA 200 100 85.0 2.06 27.4 85.0 2.07 27.5 92.0 11.7 4.8 92.0 12.8 3.7 86.0 1.8 25.7
s.e (10.0) (2.6) (21.6) (10.0) (2.6) (21.6) (2.0) (7.7) (3.2) (3.0) (7.8) (2.2) (10.0) (2.3) (21.7)
1000 250 98.3 2.7 0.7 98.0 2.7 0.7 93.0 13.2 6.3 90.5 16.7 11.8 98.5 2.4 0.6
s.e (3.0) (5.5) (1.3) (3.0) (5.5) (1.4) (12.5) (24.6) (24.3) (16.0) (31.4) (32.3) (2.4) (4.9) (1.1)
500 300 94.0 1.3 9.8 94.0 1.3 9.7 96.0 7.8 0.3 96.0 8.8 0.1 95.0 1.1 9.4
s.e (7.0) (2.4) (14.4) (7.0) (2.5) (14.2) (4.0) (8.3) (0.5) (4.0) (8.7) (0.2) (7.0) (2.1) (14.1)
CSN DATA 200 100 86.0 5.0 22.8 86.0 5.0 22.2 88.0 17.3 6.3 88.0 18.5 5.7 83.0 4.4 28.9
s.e (10.0) (4.9) (21.1) (10.0) (4.9) (20.8) (4.0) (10.2) (2.1) (5.0) (10.2) (2.1) (10.0) (4.7) (24.1)
1000 250 98.0 1.6 3.2 98.0 1.6 3.0 96.0 8.3 0.7 94.2 11.0 4.5 96.0 1.3 6.7
s.e (3.0) (2.5) (6.6) (3.0) (2.5) (6.2) (5.7) (11.3) (8.1) (10.0) (20.6) (20.7) (6.0) (2.2) (12.5)
500 300 92.0 1.5 14.7 92.1 1.6 14.2 95.0 8.3 1.7 95.0 9.4 0.9 89.5 1.2 19.7
s.e (9.1) (2.2) (19.3) (8.9) (2.2) (18.8) (3.4) (7.8) (1.6) (4.0) (8.3) (0.6) (11.7) (1.8) (24.4)
SUT DATA 200 100 80.0 2.9 37.4 80.0 3.0 36.7 (-) (-) (-) 89.0 15.6 7.1 71.0 2.2 54.9
s.e (10.0) (3.8) (24.2) (10.0) (3.9) (37.4) (-) (-) (-) (2.0) (6.3) (1.7) (12.0) (3.2) (26.8)
1000 250 97.0 2.1 4.7 97.0 2.1 4.5 (-) (-) (-) 96.0 7.7 0.1 89.0 1.7 20.3
s.e (4.0) (4.5) (8.6) (4.0) (4.5) (8.3) (-) (-) (-) (3.7) (7.4) (0.01) (12.0) (4.2) (25.2)
500 300 89.3 1.9 19.8 89.6 1.9 18.9 (-) (-) (-) 94.4 9.7 1.5 80.0 1.6 37.6
s.e (9.1) (3.7) (19.8) (9.0) (3.7) (19.5) (-) (-) (-) (3.3) (6.9) (0.7) (14.4) (3.3) (30.5)
@@ 6.2 Empirical Analysis: S&P500 Log-Returns
In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo! Finance, using the proposed elliptical skew-(S)KEPTIC estimators and comparing it with the SKEPTIC. After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations. The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the normal distribution (Cont, (2001)).
@@
As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials. We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component. Moreover, we investigate the Gaussian assumption of the asset log-returns by computing normality tests with different significance levels at 5% and 1%. Table 3 indicates a clear and complete deviation from normality in the stock data. This could be attributed to the presence of exceptional events, thus a huge presence of outlier values during the analyzed period, which lead to a departure from the assumption of normal distribution for log-returns.
@@
To estimate the graphs, given the large number of variables relative to observations we determine the tuning parameter $\lambda_{G}$ using the Extended Bayesian Information Criteria (EBIC) of Chen and Chen, (2008) to ensure a balanced trade-off between model fit and complexity.
$\operatorname{EBIC}=-2L+E\log(n)+4\gamma E\log(p).$ (19)
The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes. We set $\gamma=1$, ensuring a strong penalization, as desired. The tuning parameter $\lambda_{G}$ is set equal to $0.54$ the value that minimizes the scores of the first-order differences in EBIC. To align the log-returns to the unified skew-$t$ distribution for computing the elliptical skew-KEPTIC estimator, we estimate the degrees of freedom of the distribution. We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method. In likelihood methods, the underlying assumption is that the observations are independent, therefore, we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the R package sn developed by Azzalini, (2023).
@@
The stocks that belong to the same sector tend to cluster, and those at the periphery of the network are uncorrelated with the others during the considered period, thus, they are uncorrelated with the movement of the market. Peralta and Zareei, (2016) show that those stocks have a greater weight in constructing a portfolio using the Markowitz approach. Therefore, in large-dimensions, a network approach can be valuable for identifying stocks with the least correlation to the market.
Table 4 demonstrates that incorporating skewness and kurtosis components leads to a sparser network estimation compared to the normal case.
@@ 7 Concluding remarks
We propose rank-based correlation matrix estimators, named the elliptical skew-(S)KEPTIC estimators, to provide semiparametric estimation of graphical models built on meta skew-elliptical distributions. The proposed framework includes two classes of estimators: the elliptical skew-KEPTIC estimator which employs Kendall’s tau statistic, and the elliptical skew-SKEPTIC estimator which also uses Spearman’s rho statistic to estimate the correlation matrix.
We study the theoretical properties of the rank-based correlation matrix estimators and establish concentration bounds. These bounds yield statistical rates of convergence under the elementwise maximum norm for both the correlation matrix estimation and the resulting estimation of the precision matrix. We also study the sparsistency of the precision matrix, evaluating the performance of the elliptical skew-(S)KEPTIC estimators in correctly recovering the edges of the graphical model.
The empirical analysis and numerical simulations demonstrate that the elliptical skew-(S)KEPTIC estimators yield the optimal performance when the data deviates from normality. Applied to real data, specifically the log-returns of S&P500 constituent stocks, the resulting graphs are sparser compared to those estimated under normality.
Our findings show a reliable framework that is useful when dealing with data that exhibit high skewness which cannot be neglected, and its effectiveness is validated through the optimal graph recoveries of the numerical simulations.
@@ Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and
$Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$. We compute the correlation matrix of $\boldsymbol{X}$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$ exploiting (8).
$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}(I-D^{2})L\right)^{-1},$
$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2}.$
The matrices $I$ and $D^{2}$ are both diagonal matrices, and the matrix $M$ is. Thus, we have a new decomposition of the correlation matrix in which the skewness parameter $\alpha$, belonging to the matrix $D_{\alpha}$, contributes to the determination of the linear relations. The matrix $(L^{\top}ML)$ is:
$\left(\begin{array}[]{ccccc}m_{11}&\ell_{12}m_{11}&\ell_{13}m_{11}&\ldots&\ell_{1p}m_{11}\\ \ell_{21}m_{11}&\sum^{p-(p-2)}_{k=1}\ell_{k2}\ell_{2k}m_{kk}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{2k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}\\ \ell_{31}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{3k}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{3k}m_{kk}&\ldots&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}\\ \vdots&\ldots&\ldots&\ddots&\vdots\\ \ell_{p1}m_{11}&\sum_{k=1}^{p-(p-2)}\ell_{k2}\ell_{pk}m_{kk}&\sum_{k=1}^{p-(p-3)}\ell_{k3}\ell_{pk}m_{kk}&\ldots&\sum_{k=1}^{p}\ell_{kp}\ell_{pk}m_{kk}\end{array}\right)$
where the generic $m_{k,k}$ term is:
$m_{kk}=I_{kk}-D^{2}_{kk}=1-\frac{2}{\pi}\frac{\alpha_{kk}^{2}}{1+\alpha_{kk}^{2}}.$
Let $A_{{\tau}}\equiv L^{\top}ML$ denote the rescaled final correlation matrix of $\boldsymbol{X}$, estimated using Kendall’s tau statistic. We recall that the stochastic representation (7) of the CSN is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma,D_{\alpha}L,\nu,0,D_{\kappa})$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$. Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of the closed skew-normal distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Omega^{-1},D_{\alpha}L,\nu,0,D_{\kappa})$. ∎
@@
We recall Lemma 2.1, where for the normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$. Let $A_{\rho}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{t}{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$. ∎

### v3

@@ 6.1 Numerical Simulations
Table 2: Quantitative comparisons of the 6 correlation matrix estimators using different distributions to generate data, and compute the $\Omega$ matrix. The graphs are estimated using the glasso algorithm. The values are in percentages.
SKEPTIC ($\widehat{\rho}$) SKEPTIC ($\widehat{\tau}$) SKEW-SKEPTIC ($\widehat{\rho}$) SKEW-SKEPTIC ($\widehat{\tau}$) SKEW-KEPTIC PEARSON
AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR
r = 0.05 $\gamma$ = 1.5 91.0 6.8 11.8 91.0 6.9 11.5 91.0 6.6 12.2 91.0 6.7 11.9 90.0 6.4 13.6 54.0 14.6 78.4
s.e (1.3) (2.1) (3.5) (1.4) (2.1) (3.5) (1.4) (2.0) (3.6) (1.4) (2.0) (3.6) (1.7) (2.0) (4.4) (1.4) (11.7) (9.9)
r = 0.05 $\gamma$ = 2.5 87.8 17.8 6.5 88.8 17.9 6.4 88.8 17.6 6.8 88.8 17.7 6.6 88.7 17.5 7.2 50.1 30.9 68.3
s.e (2.7) (6.0) (3.2) (2.8) (6.0) (3.2) (2.7) (5.9) (3.2) (2.7) (6.0) (3.4) (2.6) (6.0) (3.3) (0.9) (15.4) (14.3)
r = 0.05 $\gamma$ = 3 84.8 20.8 9.5 84.8 21.0 9.4 84.8 20.7 9.7 84.8 20.8 9.6 84.7 20.6 10.0 49.6 41.5 58.1
s.e (12.3) (10.5) (14.3) (12.3) (10.5) (14.4) (12.3) (10.6) (14.3) (12.3) (10.6) (14.3) (12.3) (10.6) (14.2) (0.7) (13.4) (12.8)
r = 0.1 $\gamma$ = 1.5 81.0 3.4 34.6 81.4 3.5 33.7 80.6 3.3 35.5 81.0 3.4 34.6 80.1 3.2 36.7 50.4 24.2 74.8
s.e (3.0) (2.0) (6.9) (3.0) (2.0) (6.9) (3.0) (1.9) (6.8) (3.0) (1.9) (7.0) (3.0) (1.9) (6.8) (0.5) (17.9) (17.2)
r = 0.1 $\gamma$ = 2.5 84.8 7.8 22.8 85.1 7.7 22.2 84.6 7.4 23.4 84.8 7.5 22.9 84.5 7.3 23.8 49.5 39.8 59.9
s.e (1.8) (1.8) (4.9) (1.7) (1.8) (4.6) (1.8) (1.8) (5.0) (1.8) (1.8) (4.9) (1.8) (1.8) (4.9) (0.6) (15.6) (15.1)
r = 0.1 $\gamma$ = 3 86.6 12.8 14.0 86.7 13.0 13.6 86.4 12.7 14.4 86.6 12.8 14.0 86.3 12.6 14.8 50.0 31.6 67.9
s.e (2.2) (3.4) (6.0) (2.2) (3.4) (5.7) (2.3) (3.6) (6.1) (2.2) (3.4) (5.8) (2.3) (3.4) (6.0) (0.7) (15.0) (14.5)
r = 0.2 $\gamma$ = 1.5 68.1 2.9 61.0 68.5 2.9 60.1 67.8 2.8 61.6 68.2 2.8 60.9 67.5 2.7 62.4 50.2 31.8 67.7
s.e (4.2) (1.5) (9.5) (4.3) (1.5) (9.6) (4.2) (1.5) (9.5) (4.3) (1.5) (9.6) (4.3) (1.4) (9.6) (0.4) (12.3) (11.8)
r = 0.2 $\gamma$ = 2.5 70.5 5.3 53.7 70.8 5.4 53.0 70.2 5.2 54.4 70.5 5.3 53.7 70.3 5.2 54.3 50.1 31.1 68.4
s.e (3.1) (1.8) (7.8) (3.3) (1.8) (8.0) (3.2) (1.8) (7.9) (3.3) (1.8) (8.1) (3.3) (1.8) (8.1) (0.6) (15.6) (14.9)
r = 0.2 $\gamma$ = 3 71.9 6.6 49.6 72.3 6.6 48.8 71.7 6.4 50.2 72.0 6.5 49.5 71.5 6.3 50.7 50.1 30.9 68.5
s.e (4.0) (3.3) (9.9) (3.8) (3.4) (9.7) (3.9) (3.3) (9.8) (3.8) (3.2) (9.7) (3.9) (3.2) (9.5) (0.6) (14.5) (14.0)
@@ 6.2 Empirical Analysis: S&P500 Log-Returns
In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo! Finance, using the proposed elliptical skew-(S)KEPTIC estimators and comparing them with the SKEPTIC estimator. After cleaning the data, the dataset consists of 454 stocks, each with 1762 observations. The log-returns of the stocks in the S&P500500 index are empirically skewed and exhibit heavier tails than the normal distribution (cont2001empirical).
@@
As a motivating example, we benchmark the analysis of han2014scale, reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials. We appreciate the departure from normality reflected in a leptokurtic distribution with a relevant skewness component. Moreover, we investigate the Gaussian assumption of the asset log-returns by computing normality tests with different significance levels at 1% and 5%. Table 3 indicates a clear and complete deviation from normality in the stock data. This could be attributed to the presence of exceptional events, thus a huge presence of outlier values during the analyzed period, which lead to a departure from the assumption of normal distribution for log-returns.
@@
To estimate the graphs, given the large number of variables relative to observations, we determine the tuning parameter $\lambda_{G}$ using the StARS approach of liu2010stability. The tuning parameter $\lambda_{G}$ is set equal to $0.58$. To compute the elliptical skew-(S)KEPTIC estimators, we need to estimate the skewness parameters for the closed skew-normal distribution and for the unified skew-$t$ distribution. The skewness parameters are estimated using the same procedure as described in the numerical simulations.
@@
The stocks that belong to the same sector tend to cluster, and those at the periphery of the network are uncorrelated with the others during the considered period, thus, they are uncorrelated with the movement of the market. peralta2016network show that those stocks have a greater weight in constructing a portfolio using the Markowitz approach. Therefore, in large-dimensions, a network approach can be valuable for identifying stocks with the least correlation to the market. That is, incorporating skewness and kurtosis components leads to a sparser network estimation compared to the normal case as reported in Table 4.
@@ 7 Concluding remarks
We propose rank-based correlation matrix estimators, named the elliptical skew-(S)KEPTIC estimators, to provide semiparametric estimation of graphical models built on meta skew-elliptical distributions. The proposed framework includes two classes of estimators: the elliptical skew-KEPTIC estimator, which employs Kendall’s tau statistic, and the elliptical skew-SKEPTIC estimator, which also uses Spearman’s rho statistic to estimate the correlation matrix.
We study the theoretical properties of the rank-based correlation matrix estimators and establish concentration bounds. These bounds yield statistical rates of convergence under the elementwise maximum norm for both the correlation matrix estimation and the resulting estimation of the precision matrix.
The empirical analysis and numerical simulations demonstrate that the elliptical skew-(S)KEPTIC estimators yield the optimal performance when the data deviates from normality. The application to the log-returns of the S&P500 constituent stocks shows resulting graphs that are sparser compared to those estimated under normality.
Our findings demonstrate a reliable framework that is useful when dealing with data that exhibit high skewness which cannot be neglected, and its effectiveness is validated through the optimal graph recoveries of the numerical simulations.
@@ Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}\!\big(\Omega^{-1},\,D_{\alpha}L,\,h^{(p)},\,0,\,D_{\kappa};\,f\big)$, where $f(\boldsymbol{X})\stackrel{{\scriptstyle d}}{{=}}V_{0}^{-1/2}Z_{0}$ and $Z_{0}\sim\operatorname{CSN}_{p,p}\!\big(0,\,\Omega^{-1},\,D_{\alpha}L,\,0,\,D_{\kappa}\big)$, with $D_{\kappa}=I_{p}$ and $V_{0}=1$. Under this framework, the stochastic representation of $\boldsymbol{X}$ coincides with (8); explicitly:
$\boldsymbol{X}$ $\stackrel{{\scriptstyle d}}{{=}}L^{-1}D_{\kappa}^{-1/2}D_{\alpha}(I+D_{\alpha}^{2})^{-1/2}\,\boldsymbol{U}\;+\;L^{-1}D_{\kappa}^{-1/2}(I+D_{\alpha}^{2})^{-1/2}\,\boldsymbol{V}.$
${\boldsymbol{X}}=D_{\alpha}(I+D_{\alpha}^{2})^{-1/2}\,\boldsymbol{U}\;+\;(I+D_{\alpha}^{2})^{-1/2}\,\boldsymbol{V}.$
Here $\boldsymbol{V}\sim\operatorname{NPN}(0,\Sigma^{0},f)$, where $\Sigma^{0}=L^{-1}D_{\kappa}L$, meaning that its joint CDF is:
$F_{V}(v_{1},\ldots,v_{p})=\Phi_{\Sigma^{0}}\!\big(\,\Phi^{-1}(F_{V_{1}}(v_{1})),\,\ldots,\,\Phi^{-1}(F_{V_{p}}(v_{p}))\,\big),$
where each $F_{V_{j}}$ is estimated from normalized ranks as described in Section 2.2. Regarding the $U$ distribution:
$U_{ij}=F_{\mathrm{HN}}^{-1}(\widehat{r}_{ij})=\Phi^{-1}\!\Big(\frac{\widehat{r}_{ij}+1}{2}\Big),$
where $F_{\mathrm{HN}}^{-1}$ is the standard half-normal quantile function. By copula invariance under strictly increasing marginal transforms, $\operatorname{Copula}(U)=\operatorname{Copula}(V)$, (see e,g, embrechts2002correlation), this means that $U$ and $V$ share the same Gaussian copula dependence structure $\Sigma^{0}$. Therefore, the joint CDF of $U$ (Gaussian copula with half-normal marginals) is:
$F_{U}(u_{1},\ldots,u_{p})=\Phi_{\Sigma^{0}}\!\big(\,\Phi^{-1}(F_{\mathrm{HN}}(u_{1})),\,\ldots,\,\Phi^{-1}(F_{\mathrm{HN}}(u_{p}))\,\big),\qquad u_{i}\geq 0,$
and $\Sigma^{0}$ can be estimated with the SKEPTIC estimator, where the generic element of $\Sigma^{0}_{ij}=\sin\Big(\frac{\pi}{2}\tau_{ij}\Big)$. When $V_{0}\sim\mathrm{Gamma}(\nu/2,\nu/2)$, $\boldsymbol{X}$ reduces to a unified skew-$t$ (SUT) model and $V$ is a semiparametric $t$-copula. demarta2005t show that the identity $\Sigma^{0}_{ij}=\sin\!\big(\tfrac{\pi}{2}\tau_{ij}\big)$ still holds for $t$-copula, so the dependence structure is the same for the $t$ and half-$t$ distribution.
Let $A_{0}=D_{\alpha}(I+D_{\alpha}^{2})^{-1/2}$ and $B_{0}=(I+D_{\alpha}^{2})^{-1/2}$, and denote by $S_{\tau}:=\sin\!\big(\tfrac{\pi}{2}\,T\big)$, where $T$ is the correlation matrix estimated with Kendall’s tau statistic, the elliptical skew-KEPTIC estimator is:
$S^{\tau,\alpha}\;\equiv\;\operatorname{Cor}({\boldsymbol{X}})$ $=A_{0}\,S_{\tau}\,A_{0}\;+\;B_{0}\,S_{\tau}\,B_{0}$
$=(I+D_{\alpha}^{2})^{-1/2}\,\Big(S_{\tau}+D_{\alpha}S_{\tau}D_{\alpha}\Big)\,(I+D_{\alpha}^{2})^{-1/2}.$
In the element-wise form, for $i\neq j$, the estimator is:
$S^{\tau,\alpha}_{ij}=\sin\!\Big(\tfrac{\pi}{2}\tau_{ij}\Big)\;\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}},\quad\text{and:}\quad S^{\tau,\alpha}_{ii}=1.$
@@
We recall Lemma 2.1, where for the normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$. Let $A_{0}=D_{\alpha}(I+D_{\alpha}^{2})^{-1/2}$ and $B_{0}=(I+D_{\alpha}^{2})^{-1/2}$, and denote by $S_{\rho}:=2\sin\!\big(\tfrac{\pi}{6}\,P\big)$, where $P$ is the correlation matrix estimated using Spearman’s rho statistic. Following the proof of Theorem 3.4, the elliptical skew-SKEPTIC estimator is:
$S^{\rho,\alpha}\;\equiv\;\operatorname{Cor}({\boldsymbol{X}})$ $=A_{0}\,S_{\rho}\,A_{0}\;+\;B_{0}\,S_{\rho}\,B_{0}$
$=(I+D_{\alpha}^{2})^{-1/2}\,\Big(S_{\rho}+D_{\alpha}S_{\rho}D_{\alpha}\Big)\,(I+D_{\alpha}^{2})^{-1/2}.$
In the element-wise form, for $i\neq j$, the estimator is:
$S^{\rho,\alpha}_{ij}=2\sin\!\Big(\tfrac{\pi}{6}\rho_{ij}\Big)\;\frac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}},\qquad S^{\rho,\alpha}_{ii}=1.$